# State-action control barrier functions: Imposing safety on learning-based control with low online computational costs

### v1

@@ IV Constraint tightening and contractive CBF
Similarly, an exponential CBF $B(\cdot):\mathcal{X}\to\mathbb{R}$ is called a $\lambda$-contractive exponential CBF if there exist $\lambda>0$ and $\beta\in(0,1)$ such that
@@
The definition of contractive CBFs is introduced to endow the safe set $\mathcal{S}_{B}$ with a contractive property. It enforces an energy decay when the state is on the boundary of the safe set. In particular, if $B(\cdot)$ is a $\lambda$-contractive CBF, for any $x\in\partial\mathcal{S}_{B}$, there exists an input $u\in U$ such that $B(f(x,u))\leq-\lambda$, which means that $x^{+}=f(x,u)$ is in the interior of $\mathcal{S}_{B}$. Such a property can guarantee that any approximation of $B(\cdot)$ is still a valid CBF if the approximation error is sufficiently small.
V Approximating state-action CBF
In this section, we propose a method for approximating the state-action CBF. Inspired by HJ reachability [13] and predictive CBF [14], we propose a comprehensive framework that uses the optimal value functions of a sequence of optimization problems to implicitly represent the CBF sequence $\{B_{k}\}^{\infty}_{k=1}$. By tuning some parameters, this framework can compute standard CBFs, exponential CBFs, as well as contractive CBFs. Although the exact formulation of each $B_{k}$ is unknown, we can use sampling methods to collect data tuples $(x^{\mathrm{s}},\;u^{\mathrm{s}},\;B_{k}(f(x^{\mathrm{s}},u^{\mathrm{s}}))$, and use these tuples to learn a state-action CBF.
V-A CBF generator based on HJ reachability
@@
$B_{k}(x):=$ $\min_{\{x_{t},u_{t}\}_{t=0}^{k}}\max\left\{\max_{t\in\mathbb{I}(k-1)}\alpha^{t}(h(x_{t})+\lambda_{t}),\alpha^{k}B_{0}(x_{k})\right\}$
$\quad\mathrm{s.t.}\;x_{t+1}=f\left(x_{t},u_{t}\right),\;\;u_{t}\in U,\;t=0,1,...,k-1,$
@@
In (V-A), $\mathbb{I}(k-1)$ denotes the set $\{0,1,...,k-1\}$, $\alpha\geq 1$, and $B_{0}$ is a CBF, which, however, could have a very small safe set. We take $\alpha=1$ if $B_{0}$ satisfies (3), or let $\alpha=1/\beta$ if $B_{0}$ is an exponential CBF. To solve (V-A), we need to know the explicit formulation of $B_{0}$. An optimization-based method that can compute a local quadratic $B_{0}$ for the nonlinear system (1) is reported in [15], by linearizing the model around the origin and solving an LMI problem. In Appendix -D, this method is extended to our situations where additional conditions such as (8), (9), and (11)-(12) in the subsequent Theorem 1 are required. Besides, $\lambda_{t}\geq 0,\;t=0,1,...,k-1$ are tuning parameters. We consider the following three options for choosing $\lambda_{t}$:
@@
Option 1 means that we are constructing CBFs for the system (1) under the original state constraints. If Option 2 is chosen, it is seen from (V-A) that we increase the function $h$ to $h+\lambda$, i.e., we tighten the original state constraints to $x\in X_{\lambda}=\{x\in\mathcal{X}|h(x)+\lambda\leq 0\}$. More conservatively, in Option 3, we require a linear decrease rate for $h$ w.r.t. the time step to construct contractive CBFs.
@@
Consider $B_{k}$ from (V-A). Suppose that $B_{0}$ is a (exponential) CBF for system (1) with the state constraint $x\in X$. We have the following results:
@@
    Let $\lambda_{t}=0,\;\forall t\in\{0,1,...,k-1\}$. Then $B_{k},\;k=1,2,...$ is a (exponential) CBF for system (1) with the state constraint $x\in X$.
@@
    Let $\lambda_{t}=\lambda>0,\;\forall t\in\{0,1,...,k-1\}$. If $\lambda,B_{0}$ satisfy
@@
    Let $\lambda_{t}=t\lambda,\;\forall t\in\{0,1,...,k-1\}$, $\lambda>0$. If $B_{0}$ is a $\lambda$-contractive (exponential) CBF and $k,\lambda,B_{0}$ satisfy
@@
    In the statements 1-3, $B_{k}(\cdot)$ is a continuous function in $\mathcal{X}$, and the safe sets satisfy $\mathcal{S}_{B_{0}}\subseteq\mathcal{S}_{B_{1}}\subseteq...\subseteq\mathcal{S}_{B_{\infty}}$. Furthermore, if $f$, $h$, and $B_{0}$ are Lipschitz continuous, $B_{k}$ is Lipschitz continuous.
@@
The condition (12) indicates that the horizon $k$ should not be selected too large when using (V-A) to generate a contractive CBF.
@@
V-B Learning state-action CBF from CBF samples
After the CBF generator is provided with a fixed $k$, a regression model can be trained to obtain an approximation of a state-action CBF.
@@ V-B Learning state-action CBF from CBF samples
After $N$ samples $\{x^{\mathrm{s}}_{i},u^{\mathrm{s}}_{i}\}^{N}_{i=1}$ are collected, problem (V-A) is solved with $x$ specified as each $f(x^{\mathrm{s}}_{i},u^{\mathrm{s}}_{i})$. As a result, $N$ data tuples $\{(x^{\mathrm{s}}_{i},\;u^{\mathrm{s}}_{i},\;Q(x^{\mathrm{s}}_{i},u^{\mathrm{s}}_{i}))\}^{N}_{i=1}$ are obtained after substituting $x=x^{\mathrm{s}}_{i},\;u=u^{\mathrm{s}}_{i}$ into (13). For general nonlinear systems, (V-A) is usually a nonlinear non-convex optimization problem, which requires a multi-start strategy [23] to find a sufficiently good optimum. For linear systems with linear or ellipsoidal constraints, (V-A) is a convex quadratically constrained quadratic program (QCQP) and the global optimum can be conveniently obtained via gradient-based optimization methods. For piecewise affine systems with piecewise affine constraints, (V-A) can be regarded as a mixed-integer QCQP, which can be globally solved by the branch-and-bound approach [24].
In this work, we use neural networks (NNs) to approximate the state-action CBF. For the parameterization (7), we use three NNs to represent $q_{1,\theta},\;q_{2,\theta}$ and $Q_{3,\theta}$ in (7), respectively. There are several ways to guarantee the positive-semidefiniteness of $Q_{3,\theta}$. For example, we can further parameterize $Q_{3,\theta}$ by $Q_{3,\theta}=L_{\theta}L^{T}_{\theta}$, where $L_{\theta}\in\mathbb{R}^{n_{u}\times n_{u}}$ is a lower triangular matrix with non-negative diagonal entries. Alternatively, we can parameterize $Q_{3,\theta}$ by $Q_{3,\theta}=P_{0}\mathrm{diag}(r_{\theta})\mathrm{diag}(r_{\theta})P^{T}_{0}$, where $r_{\theta}\in\mathbb{R}^{n_{u}}$ is the output of an NN and $P_{0}\in\mathbb{R}^{n_{u}\times n_{u}}$ is a given matrix.
With the specified regression model and the training data, the parameter $\theta$ is optimized to minimize the mean square error of $Q$ over all state-action samples.
@@ V-C Performance analysis
After the approximation $Q_{\theta}$ is obtained, it can be integrated into the safety filter (6), i.e., $Q$ in (6) is replaced by $Q_{\theta}$. In general, the safety of the policy generated from (6) will not be ensured due to the approximation error. However, in this section, we will demonstrate that we can guarantee safety in the presence of a small approximation error. To achieve this, we need an assumption on the boundedness of this error.
@@
$|Q_{\theta}(x,u)-Q(x,u)|\leq\Delta,\;\forall(x,u)\in\Omega\times U.$ (14)
Assumption 1 is common in the literature studying performance guarantees of learning-based control [25, 26, 14, 27]. If the chosen function approximator represents a continuous function, since $Q$ is also continuous, the approximation error is always upper bounded in any compact region. To obtain the upper bound, we can first get the error bound on a finite number of samples, and then extract a statistical estimate [25] or compute a deterministic bound in the whole region by using the Lipschitz property of $Q$ and $Q_{\theta}$ [27].
@@
$\pi_{\theta}(x)=\arg\min\{||u-\pi_{0}(x)||_{2},\;\mathrm{s.t.}\;u\in U,\;Q_{\theta}(x,u)\leq\Delta_{i}\},$ (15)
where $i=1,2$, or $3$, corresponding to Options 1,2, and 3 chosen in the CBF generator (V-A), $\Delta_{1}=\Delta_{2}=0$ and $\Delta_{3}=-\lambda+\Delta$.
@@
Theorem 2
@@
Consider the system (1) controlled by the policy $\pi_{\theta}$, the state-action CBF $Q$ from (13), and the safety filter (15). Suppose that Assumption 1 is fulfilled. We have the following results: (i) If $B_{k}$ is a CBF for the original state constraint, for any initial state $x_{0}\in\mathcal{S}_{B_{k}}$ that makes problem (15) recursively feasible, the closed-loop system $x_{t+1}=f\left(x_{t},\pi_{\theta}(x_{t})\right),\;t=0,1,\ldots$ has the maximum constraint violation $\Delta$, i.e., $\max_{t\in\mathbb{I}(\infty)}\{h(x_{t})\}\leq\Delta$. (ii) If $B_{k}$ is a CBF for the tightened state constraint $x\in X_{\lambda}$ and $\Delta\leq\lambda$, for any initial state $x_{0}\in\mathcal{S}_{B_{k}}$ that makes problem (15) recursively feasible, the closed-loop system $x_{t+1}=f\left(x_{t},\pi_{\theta}(x_{t})\right)$ always satisfies the state constraint $x\in X$, i.e., $\max_{t\in\mathbb{I}(\infty)}\{h(x_{t})\}\leq 0$. (iii) If $B_{k}$ is a $\lambda$-contractive CBF for the original state constraint and $\Delta\leq\lambda/2$, problem (15) will be recursively feasible for all initial states in $\mathcal{S}_{B_{k}}$ and the closed-loop system $x_{t+1}=f\left(x_{t},\pi_{\theta}(x_{t})\right)$ always satisfies the state constraint $x\in X$, i.e., $\max_{t\in\mathbb{I}(\infty)}\{h(x_{t})\}\leq 0$.
@@
In other words, in Option 2 of (V-A), we get a CBF $B_{k}$ for the tightened constraint $x\in X_{\lambda}$. The policy $\pi_{\theta}$ will make the system always satisfy the original state constraint if problem (15) is always feasible. However, the feasibility of problem (15) is only guaranteed for the initial state, i.e., problem (15) can be infeasible for some subsequent states. Practically, ones need to perform offline some statistical or deterministic verification [25] to analyze the recursive feasibility of (15). Otherwise, as recursive feasibility of (15) cannot be predicted in advance, one needs to do some relaxation for the constraint in (15) once infeasibility is observed. In comparison, in Option 3 of (V-A), we get a $\lambda$-contractive CBF $B_{k}$. An induced feature is the recursive feasibility of problem (15). As a result, under the sufficiently small approximation error, the state-action CBF approximation $Q_{\theta}$ will render $\pi_{\theta}$ a safe policy according to Definition 2.
V-D Discussion on state-action CBFs

### v2

@@ IV Constraint tightening and contractive-set CBFs
Similarly, an exponential CBF $B(\cdot):\mathcal{X}\to\mathbb{R}$ is called a $\lambda$-contractive-set exponential CBF if there exist $\lambda>0$ and $\beta\in(0,1)$ such that
@@
The definition of contractive-set CBFs is introduced to endow the safe set $\mathcal{S}_{B}$ with a contractive property. It enforces an energy decay when the state is on the boundary of the safe set. In particular, if $B(\cdot)$ is a $\lambda$-contractive-set CBF, for any $x\in\partial\mathcal{S}_{B}$, there exists an input $u\in U$ such that $B(f(x,u))\leq-\lambda$, which means that $x^{+}=f(x,u)$ is in the interior of $\mathcal{S}_{B}$. Such a property can guarantee that any approximation of $B(\cdot)$ is still a valid CBF if the approximation error is sufficiently small.
V Approximating SACBFs
In this section, we propose a method for approximating the SACBF. Inspired by HJ reachability [19] and predictive CBF [13], we propose a comprehensive framework that uses the optimal value functions of a sequence of optimization problems to implicitly represent the CBF sequence $\{B_{k}\}^{\infty}_{k=1}$. By tuning some parameters, this framework can compute standard CBFs, exponential CBFs, as well as contractive-set CBFs. Although the exact expression of each $B_{k}$ is in general intractable to compute, we can use learning-based approaches such as supervised learning and RL [25] and then use the results of Lemma 1 to approximate SACBFs.
V-A CBF generator based on reachability analysis
@@
$B_{k}(x):=$ $\min_{\{x_{t},u_{t}\}_{t=0}^{k}}\max\left\{\max_{t\in\mathcal{I}(k-1)}\alpha^{t}(h(x_{t})+\lambda_{t}),\alpha^{k}B_{0}(x_{k})\right\}$
$\quad\mathrm{s.t.}\;x_{t+1}=f\left(x_{t},u_{t}\right),\;\;u_{t}\in U,\;t\in\mathcal{I}(k-1),$
@@
In (V-A), $\mathcal{I}(k-1)=\{0,1,...,k-1\}$, $\alpha\geq 1$, and $B_{0}$ is a CBF, which, however, could have a very small safe set. We take $\alpha=1$ if the CBF $B_{0}$ is not an exponential CBF, or $\alpha=1/\beta$ if $B_{0}$ is an exponential CBF. To solve (V-A), we need to know the explicit formulation of $B_{0}$. An LMI-based method that can compute a local quadratic $B_{0}$ for the nonlinear system (1) is reported in [20]. In Appendix -D, this method is extended to our situations where additional conditions such as (8), (9), and (11)-(12) in the subsequent Theorem 1 are required. The LMI method to obtain $B_{0}$ relies on the linearization of both the system and the constraints. The resulting $B_{0}$ is in general very conservative. In contrast, by introducing the reachability problem (V-A), we get less conservative CBFs $B_{k}$, as stated in Theorem 1 below.
In (V-A), $\lambda_{t}\geq 0,\;t\in\mathcal{I}(k-1)$ are tuning parameters. We consider the following three options for choosing $\lambda_{t}$:
@@
Option 1 means that we are constructing CBFs for the system (1) under the original state constraints. If Option 2 is chosen, it is seen from (V-A) that we increase the function $h$ to $h+\lambda$, i.e., we tighten the original state constraints to $x\in X_{\lambda}=\{x\in\mathcal{X}|h(x)+\lambda\leq 0\}$. More conservatively, in Option 3, we require a linear decrease rate for $h$ w.r.t. the time step to construct contractive-set CBFs.
@@
Consider $B_{k}$ from (V-A). Suppose that $B_{0}$ is a (exponential) CBF for (1) with the state constraint $x\in X$.
@@
    Let $\lambda_{t}=0,\;t\in\mathcal{I}(k-1)$. Then $B_{k},\;k=1,2,...$ is a (exponential) CBF for system (1) with the state constraint $x\in X$.
@@
    Let $\lambda_{t}=\lambda>0,\;t\in\mathcal{I}(k-1)$. If $\lambda,B_{0}$ satisfy
@@
    Let $\lambda_{t}=t\lambda,\;t\in\mathcal{I}(k-1)$, $\lambda>0$. If $B_{0}$ is a $\lambda$-contractive (exponential) CBF and $k,\lambda,B_{0}$ satisfy
@@
    In the statements (i)-(iii), $B_{k}(\cdot)$ is a continuous function in $\mathcal{X}$, and the safe sets satisfy $\mathcal{S}_{B_{0}}\subseteq\mathcal{S}_{B_{1}}\subseteq...\subseteq\mathcal{S}_{B_{\infty}}$. Furthermore, if $f$, $h$, and $B_{0}$ are Lipschitz continuous, $B_{k}$ is Lipschitz continuous.
@@
The condition (12) indicates that the horizon $k$ should not be selected too large when using (V-A) to generate a contractive-set CBF.
@@
V-B Learning SACBFs from CBF samples
We assume that the prediction model in the reachability problem (V-A) is known. To approximate SACBFs, we opt for supervised learning, which is more computationally efficient than RL because it can solve (V-A) exactly and efficiently for each state sample.
After the CBF generator is provided with a fixed $k$, a regression model can be trained to obtain an approximation of an SACBF.
@@ V-B Learning SACBFs from CBF samples
After $N$ samples $\{x^{(\mathrm{s})}_{i},u^{(\mathrm{s})}_{i}\}^{N}_{i=1}$ are collected, problem (V-A) is solved with $x$ specified as each $f(x^{(\mathrm{s})}_{i},u^{(\mathrm{s})}_{i})$. As a result, $N$ data tuples $\{(x^{(\mathrm{s})}_{i},\;u^{(\mathrm{s})}_{i},\;Q(x^{(\mathrm{s})}_{i},u^{(\mathrm{s})}_{i}))\}^{N}_{i=1}$ are obtained after substituting $x=x^{(\mathrm{s})}_{i},\;u=u^{(\mathrm{s})}_{i}$ into (13). For general nonlinear systems, (V-A) is usually a nonlinear non-convex optimization problem, which requires a multi-start strategy [31] to find a sufficiently good optimum. For linear systems with linear or ellipsoidal constraints, (V-A) is a convex quadratically constrained quadratic program (QCQP) and the global optimum can be conveniently obtained via gradient-based optimization methods. For piecewise affine systems with piecewise affine constraints, (V-A) can be regarded as a mixed-integer QCQP, which can be globally solved by the branch-and-bound approach [32].
In this work, we use NNs to approximate the SACBF. For the parameterization (7), we use three NNs to represent $q_{1,\theta},\;q_{2,\theta}$, and $Q_{3,\theta}$ in (7), respectively. There are several ways to guarantee the positive-semidefiniteness of $Q_{3,\theta}$. For example, we can further parameterize $Q_{3,\theta}$ by $Q_{3,\theta}=L_{\theta}L^{T}_{\theta}$, where $L_{\theta}\in\mathbb{R}^{n_{u}\times n_{u}}$ is a lower triangular matrix with non-negative diagonal entries. Alternatively, we can parameterize $Q_{3,\theta}$ by $Q_{3,\theta}=P_{0}\mathrm{diag}(r_{\theta})\mathrm{diag}(r_{\theta})P^{T}_{0}$, where $r_{\theta}\in\mathbb{R}^{n_{u}}$ is the output of an NN and $P_{0}\in\mathbb{R}^{n_{u}\times n_{u}}$ is a given matrix.
Finally, $\theta$ is optimized to minimize the mean square error between $B_{k}(f)$ and $Q_{\theta}$ over all state-action samples, using classical NN training algorithms such as stochastic gradient descent [33].
@@ V-C Performance analysis
After the approximation $Q_{\theta}$ has been obtained, it can be integrated into the safety filter (6), i.e., $Q$ in (6) is replaced by $Q_{\theta}$. In general, the safety of the policy generated from (6) will not be ensured due to the approximation error. However, we will now demonstrate that we can guarantee safety in the presence of a small approximation error. To achieve this, we need an assumption on the boundedness of this error.
@@
$|Q_{\theta}(x,u)-Q(x,u)|\leq\Delta,\;\forall(x,u)\in\Omega\times U,$ (14)
where $\Omega$ is a compact set and satisfies $\mathcal{S}_{Q}\subseteq\Omega$.
Assumption 1 is common in the literature studying performance guarantees of learning-based control [12, 13, 34]. If the chosen function approximator represents a continuous function, since $Q$ is also continuous, the approximation error is always upper bounded in any compact region. To obtain the upper bound, we can first get the error bound on a finite number of samples, and then extract a statistical estimate [12] or compute a deterministic bound in the whole region by using the Lipschitz property of $Q$ and $Q_{\theta}$ [34].
@@
$\pi_{\theta}(x)=\arg\min_{u\in U}$ $||u-\pi_{0}(x)||_{2}$
$\mathrm{s.t.}\;$ $Q_{\theta}(x,u)\leq 0\text{ for Options 1 and 2 }$
$Q_{\theta}(x,u)\leq-\lambda+\Delta\text{ for Option 3}.$ (15)
@@
Theorem 2 (Safety and recursive feasibility)
Consider the system (1) controlled by $\pi_{\theta}$, the SACBF $Q$ from (13), and the safety filter (V-C). Suppose that Assumption 1 holds.
@@
(i) If $B_{k}$ is a CBF for the original state constraint, for any initial state $x_{0}\in\mathcal{S}_{B_{k}}$ that makes problem (V-C) recursively feasible, the closed-loop system $x_{t+1}=f\left(x_{t},\pi_{\theta}(x_{t})\right),\;t=0,1,\ldots$ has the maximum constraint violation $\Delta$, i.e., $\max_{t\in\mathcal{I}(\infty)}\{h(x_{t})\}\leq\Delta$.
(ii) If $B_{k}$ is a CBF for the tightened state constraint $x\in X_{\lambda}$ and $\Delta\leq\lambda$, for any initial state $x_{0}\in\mathcal{S}_{B_{k}}$ that makes problem (V-C) recursively feasible, the closed-loop system $x_{t+1}=f\left(x_{t},\pi_{\theta}(x_{t})\right)$ always satisfies the state constraint $x\in X$, i.e., $\max_{t\in\mathcal{I}(\infty)}\{h(x_{t})\}\leq 0$.
(iii) If $B_{k}$ is a $\lambda$-contractive-set CBF for the original state constraint and $\Delta\leq\lambda/2$, problem (V-C) will be recursively feasible for all initial states in $\mathcal{S}_{B_{k}}$ and the closed-loop system $x_{t+1}=f\left(x_{t},\pi_{\theta}(x_{t})\right)$ always satisfies the state constraint $x\in X$, i.e., $\max_{t\in\mathcal{I}(\infty)}\{h(x_{t})\}\leq 0$.
@@
In other words, in Option 2 of (V-A), we get a CBF $B_{k}$ for the tightened constraint $x\in X_{\lambda}$. The policy $\pi_{\theta}$ will always make the system satisfy the original state constraint if (V-C) is always feasible. However, the feasibility of problem (V-C) is only guaranteed for the initial state, i.e., problem (V-C) can be infeasible for some subsequent states. Practically, ones need to perform offline some statistical or deterministic verification [12] to analyze the recursive feasibility of (V-C). Otherwise, as recursive feasibility of (V-C) cannot be predicted in advance, one needs to do some relaxation for the constraint in (V-C) once infeasibility is observed. In comparison, in Option 3 of (V-A), we get a $\lambda$-contractive-set CBF $B_{k}$. An induced feature is the recursive feasibility of problem (V-C). As a result, under the sufficiently small approximation error, the SACBF approximation $Q_{\theta}$ will render $\pi_{\theta}$ a safe policy according to Definition 2.
The proposed constraint tightening approach naturally has robustness to model uncertainty. In particular, suppose that (1) is a nominal system, and the real system is $x_{t+1}=f_{\mathrm{r}}(x_{t},u_{t})$. The mismatch between $f$ and $f_{\mathrm{r}}$ can be handled similarly to the learning error of $Q$. Formally, we have the following corollary:
Corollary 1 (Robustness to model uncertainty)
Suppose that Assumption 1 is fulfilled and the nominal system satisfies $\left|f(x,u)-f_{\mathrm{r}}(x,u)\right|\leq\kappa,\forall(x,u)\in\Omega\times U$, where $\kappa\geq 0$. Then, if $B_{k}$ is a $\lambda$-contractive-set CBF for the nominal system under the original state constraint and $\lambda\geq 2\Delta+\mathrm{L}_{B}\kappa$ with $\mathrm{L}_{B}$ the Lipschitz constant of $B_{k}$, problem (21) will be recursively feasible for the real system with all initial states in $\mathcal{S}_{B_{k}}$ and the closed-loop system $x_{t+1}=f_{\mathrm{r}}\left(x_{t},\pi_{\theta}(x_{t})\right)$ always satisfies the state constraint $x\in X$, i.e., $\max_{t\in\mathcal{I}(\infty)}\{h(x_{t})\}\leq 0$.
The proof is similar to the proof of (iii) of Theorem 2 and is thus omitted here.
V-D Discussion on SACBFs
In addition to the computational benefit bought by introducing SACBFs, another advantage is that it enables the synthesis of both safe controllers and safe certificates even when the explicit form of the model is unknown. The similarity between SACBFs and Q value functions in RL indicates that it is possible to use RL methods [35] such as Q value iteration or Q-learning to learn SACBFs directly without knowing the explicit model. It should be noticed that even though there are several data-driven methods to synthesize standard CBFs from black-box models [17, 36], a prediction model is necessary to implement the safety filter (5). In comparison, the proposed SACBF provides the possibility of bridging the gap between model-free learning control and CBF-based safety filters. This technological roadmap, which we refer to as direct data-driven safe control and plan to explore in future work, has been recognized as a promising direction in recent survey papers [2, 9].